20 ml of O2 contracts to 17 ml when subjected to silent electric discharge in an ozoniser. What further contraction
would be observed if the partially ozonised oxygen is treated with oil of cinnamon (which absorbs ozone). All
volumes are reported at the same temperature and pressure

Grade:12th pass

1 Answers

Sunil Kumar FP
9 years ago
We have
3O2 ----2O3
at constant temperature and pressure number of mole is directly proportional to volume
3 mole give 2 mole of ozone
20 ml contracted to 17 ml
volume of ozone remained=3 ml(20-17)
volume of O3 formed from O2 is 2/3*17 ml =11.33 ml
volume of O2 left=17-11.33 ml=5.67ml